MONT 106Q – Mathematical Thinking

Study/Discussion Questions on the curious incident of the dog in the night-time^{1} August/September, 2016

0. Playing the role of “armchair psychiatrist,” do you know a likely candidate for the clinical diagnosis of Christopher’s mental disability? Do you know anyone who has a similar condition?

1. On pages 45 - 48, Christopher describes his “Behavioral Problems” and the effect they have had on his parents and their marriage. What is the effect on the reader of the dispassionate style in which he relates this information? What challenges does this book present to the ways we usually think about characters in novels? For instance, is Christopher motivated by the sort of emotions of love or desire we usually see in fictional characters? How does he describe what it means to him for a parent to love a child? Is there any evidence that Christopher feels what we would call a normal sense of attachment to other people at all – to his parents? to his teacher Siobhan? To what does he feel a sense of attachment?

2. On page 2, Christopher says he likes the idea of a world with no people in it. Later he dreams of being an astronaut, alone in space (pages 50-51), and later still that a virus has killed off everyone except him or people like him (pages 198 - 200). What are the aspects of his make-up that make dealing with other people so difficult for him? What aspects of “normal” people’s lives and behavior is he simply unable to understand?

3. But on the other hand, do Christopher’s parents really seem to understand themselves or other people either? For instance, do they seem to understand the reasons she has been unfaithful to her husband, the reasons he invented the story about her illness and death and hid her letters from Christopher, the reasons he killed the dog Wellington, why they both seem so angry and out of control with each other much of the time? Are they (or we) really so different from Christopher? Are they (or we) just better at making guesses about what others are feeling?

4. Why are aspects of everyday city life (like dealing with buying a train ticket, getting on a subway car, navigating streets, etc.) that most “normal people” would deal with easily so difficult for Christopher? What does he say about the way he is finally able to find the train station on his trip to London?

5. One might say that Christopher experiences the world quantitatively and logically rather than emotionally. What does Christopher’s teacher Mr. Jeavons say about the reason mathematics is attractive to him? But does Christopher agree with that? What does he say about this when he describes the way he figured out the “Monty Hall Problem.” Does Mr. Jeavons underestimate the complexity of Christopher’s intellect and his responses to mathematical ideas? Does Siobhan understand Christopher better?

6. In his book, An Anthropologist on Mars, the neurologist Oliver Sacks quotes Dr. Hans Asperger as saying that people like Christopher “have a sort of intelligence scarcely

1 Adapted from www.readinggroupguides.com/reviews/the-curious-incident-of-
the-dog-in-the-night-time/guide

touched by tradition or culture – unconventional, unorthodox, strangely pure and original, akin to the intelligence of pure creativity.” Does the novel’s intensive look at Christopher’s fascinating and often profound mental life suggest that the the pity that well-meaning, “normal” people might feel for him is misdirected? Given his mathematical gifts, does his future as a mathematician or scientist look promising?

7. Is there a difference between solving a mathematical problem and navigating through life? If so, how are they different? How does the Monte Hall Problem illustrate Christopher’s perspective on that question? Is there a problem with the way most people lead their lives?

8. The author of the novel, Mark Haddon, has said about his own experience writing the book that “it’s not just a book about disability ... [Christopher is] a character [who] if you met him in real life you’d never, never get inside his head. Yet something magical happens when you write a novel about him. You slip inside his head and it seems like the most natural thing in the world.” On the other hand the author Jay McInerney, in a review of The Curious Incident of the Dog in the Night-time, wrote that from his point of view as a reader of the novel, “ ... the gulf between Christopher and his parents, between Christopher and the rest of us, remains immense and mysterious ... Christopher Boone is an unsolved mystery.” Do you feel you have shared Haddon’s experience of coming to understand a truly different sort of mind? How? Or does McInerney’s sense of mystery do a better job of describing your own reactions? Is it even possible for “normal” people to understand and empathize with the experience of someone whose mind works so differently? How does this relate to our CHQ theme?